What is the equation of a line that passes thru the point (2,11) and is parrallel to x-4y=6
Found 2 solutions by Alan3354, harpazo:
Answer by Alan3354(69443)

What is the equation of a line that passes thru the point (2,11) and is parrallel to x-4y=6
----------
Find the slope, m of the line.
(y-11) = m*(x-2)

Answer by harpazo(655)

x-4y= 6
-4y=-x+6
y = (-x+6)/-4
y = (1/4)x -6/4
The slope is (1/4). The equation you are looking for has the same slope because the line is parallel.
Plug the point and the slope (1/4) into the point-slope formula. Solve the equation for y and your answer will be the equation that is parallel to the given equation.
Take it from here.